# Inference-Time Selective Debiasing

### v1

@@ Abstract
We propose selective debiasing – an inference-time safety mechanism that aims to increase the overall quality of models in terms of prediction performance and fairness in the situation when re-training a model is prohibitive. The method is inspired by selective prediction, where some predictions that are considered low quality are discarded at inference time. In our approach, we identify the potentially biased model predictions and, instead of discarding them, we debias them using LEACE – a post-processing debiasing method. To select problematic predictions, we propose a bias quantification approach based on KL divergence, which achieves better results than standard UQ methods. Experiments with text classification datasets demonstrate that selective debiasing helps to close the performance gap between post-processing methods and at-training and pre-processing debiasing techniques.
Inference-Time Selective Debiasing
Gleb Kuzmin3,5   Nemeesh Yadav4   Ivan Smirnov5 Timothy Baldwin1,2   Artem Shelmanov1 1MBZUAI    2The University of Melbourne    3AIRI 4IIIT Delhi    5FRC CSC RAS kuzmin@airi.net    neemesh20529@iiitd.ac.in    ivs@isa.ru {timothy.baldwin, artem.shelmanov}@mbzuai.ac.ae
@@ 1 Introduction
Fairness is an important safety characteristic of a machine learning (ML) model, representing the model’s ability to classify instances without discrimination based on various sensitive attributes, such as race, gender, and age Blodgett et al. (2020). In recent years, numerous works have investigated and promoted fairness, and a variety of fairness definitions have been proposed Blodgett et al. (2020); Han et al. (2022b). One prominent type of fairness is group fairness, also known as the equal opportunity criterion, which reflects the inequality of opportunities across different groups Han et al. (2022a). The inequality in the model predictions usually comes from inadequate or biased training data, and to address this problem and achieve better fairness, researchers have proposed various debiasing techniques Li et al. (2018); Han et al. (2021, 2022a); Belrose et al. (2023). The majority of these techniques assume that one has access to the complete training data and the ability to retrain the model from scratch using some special loss function or reweighting the training instances. However, there are many situations when this assumption does not hold. There is a need for inference-time safety mechanisms that protect users from inadequate model behavior.
Inference time safety mechanisms are primarily associated with uncertainty quantification (UQ) techniques Gal and Ghahramani (2016) and selective classification Geifman and El-Yaniv (2017). Selective classification aims to enhance the reliability of ML-based applications by abstaining from unreliable predictions with high uncertainty. We suggest that the same approach could equally be applied to increase fairness.
In this work, we propose an inference-time safety mechanism that aims to increase the overall quality of models in terms of prediction performance and fairness in situations when model re-training is prohibitive. We call this approach selective debiasing. Instead of rejecting predictions of selected instances as in selective classification, we apply to them inference-time debiasing using post-processing debiasing techniques. We note that applying debiasing to all model decisions might reduce its overall performance. However, the selected predictions often have a high potential to be of low quality or carry some bias. Therefore, modifying these predictions is less risky in terms of worsening the situation, while still holding the potential to correct errors. To the best of our knowledge, this style of approach is novel to the NLP community.
@@ 2 Background
Debiasing techniques can be categorized into three groups: at-training, pre-processing, and post-processing Han et al. (2022b). The first two groups require retraining the model from scratch and access to the whole training set. They also cannot be selectively applied to a subset of predictions. Post-processing techniques do not involve changes to the model itself, can be trained on a subset of data, and can be applied to predictions selectively. However, their performance is usually worse.
@@
Adversarial training (Adv) Li et al. (2018) aims to solve a minimax game between minimizing the loss for the primary task and maximizing the loss for predicting the protected attribute. Diverse adversaries (DAdv) Han et al. (2021) extends Adv by using an ensemble of diverse discriminators instead of just one. Balanced Training with Equal Opportunity (BTEO) Han et al. (2022a) is a dataset balancing technique. It aims to minimize the TPR gap between two protected groups. Balanced Training with Joint balance (BTJ) Lahoti et al. (2020) aims to improve the worst-case performance over all unobserved protected groups by focusing on the computationally identifiable regions of error.
@@
INLP proposes an iterative method that involves finding an orthogonal projection of a linear classifier matrix, which is initially learned to predict protected attributes from representations (e.g. hidden states of the standard model). This orthogonal projection is then iteratively used to remove all relevant information from these representations, which was used by the classifier to predict protected attributes.
LEACE is a concept erasure technique that renders representations impervious to the prediction of a specific concept while minimizing change to the original representations. LEACE first whitens the data by equalizing the variance in all directions. Then the data is orthogonally projected onto the relevant subspace, and finally, the data is unwhitened with the same covariance matrix before subtracting from the original representation.
3 Proposed Methods
@@
Selective classification is a well-known human-in-the-loop approach that makes predictions for some instances and dispatches others to human experts for labeling. Instead of rejecting instances completely as in selective classification, we suggest applying debiasing to selected model predictions during the inference time. This is different from the standard post-processing debiasing methods since we suggest changing the predictions for only some instances. In particular, we identify the potentially most biased instances by a scoring criterion and replace the original model’s prediction with predictions debiased using a post-processing method. This does not introduce much computational overhead, as applying a post-processing method to a model is a matter of one or two matrix multiplications. Finally, we obtain predictions for the model that represent a mixture of original $p^{i}$ and debiased predictions $q^{i}$:
$r^{i}=\begin{cases}p^{i},\text{ if $i\in S_{h}$}\\ q^{i},\text{ if $i\notin S_{h}$},\end{cases}$ (1)
where $S_{h}$ is the set of samples with high bias scores. Membership in $S_{h}$ is determined by a threshold criterion (e.g. if the bias score is in the top-$N$ percent of highest scores). Such an approach allows flexible adjustment of the accuracy–fairness trade-off.
@@
For selective debiasing, we need a good instance-level bias quantification method. Selective classification is usually based on UQ methods. However, uncertainty on its own is not reflective of group fairness; it simply tries to eliminate potentially erroneous predictions. Figure 1 presents a motivational example. It shows the rejection plots for oracle rejection strategies in selective classification for both accuracy and fairness (see the exact definition of fairness in Appendix E). We can see that the fairness oracle outperforms the UQ oracle in terms of fairness while keeping the same performance in terms of accuracy. These results illustrate that to improve fairness without penalty to accuracy, it is necessary to change the order of instances being eliminated, i.e. we need a different selection criterion.
Consider a model that is trained to label instances with classes $c\in C$. To quantify how biased a prediction is for a given instance, we suggest using the Kullback-Leibler divergence (KL) Kullback and Leibler (1951) between the standard model prediction $p_{c}^{i}$ and the debiased prediction $q_{c}^{i}$ via LEACE for the $i$-th instance:
$U_{KL}^{i}=\sum_{c\in C}p_{c}^{i}\log\left(\frac{p_{c}^{i}}{q_{c}^{i}}\right).$ (2)
The key idea behind the usage of KL divergence is that the distance quantifies the difference in predictions between the standard and debiased model. The bigger the difference, the bigger the amount of information about the protected attribute removed from the original representation of the instance. The proposed approach allows for improvement over the baseline debiasing method by updating only the most biased instances. It is especially beneficial in the case of LEACE because it prevents excessive change in less-biased instances, which could lead to a drop in both accuracy and fairness. Moreover, selective debiasing adds a mechanism for tuning the accuracy–fairness trade-off for LEACE. This proposed method also could be applied in combination with INLP.
@@ 4 Experiments
Figure 1: Rejection results for fairness and accuracy with oracle scores on a synthetic dataset with LogReg model; the FR-AUC and Acc-AUC are the areas under fairness– and accuracy–rejection curves correspondingly. The details are presented in Appendix F.
@@ 4.1 Experimental Setup
For our experiments, we use two English text classification datasets that in addition to target variables, provide explicit protected attributes. The first is MOJI Blodgett et al. (2016), a dataset for sentiment analysis with a binary class (happy and sad) and a binary protected attribute, which corresponds to the author’s ethnicity (African American English (AAE) vs. Standard American English (SAE)). The second is a version of the widely used BIOS dataset De-Arteaga et al. (2019) for occupation classification with binary gender as the protected attribute. BIOS-2 Subramanian et al. (2021) is a two-class subsample of the original BIOS dataset with a highly-skewed joint distribution of classes and protected attribute values. As it has been shown to be beneficial to report results for both “balanced” and “imbalanced” versions of datasets Kuzmin et al. (2023), we conduct experiments on both versions. Detailed information and statistics of the datasets are presented in Appendix D.
Metrics.
We employ several metrics to quantify the predictive performance of the model and also its fairness. To estimate the model’s performance, we use accuracy. For fairness, we consider the widely-used equal opportunity criterion Hardt et al. (2016); Han et al. (2022a, b). We also use two aggregated metrics to obtain the overall model performance in terms of accuracy and fairness. The first one is the distance to the optimal point (DTO) Han et al. (2021):
$\text{DTO}=\sqrt{(1-\text{Accuracy})^{2}+(1-\text{Fairness})^{2}}.$ (3)
[…]
@@
INLP
INLP+SR,
opt. perc.
INLP+KL,
opt. perc.
MOJI imbalanced Fairness $\uparrow$ 61.8$\pm$0.7 73.7$\pm$0.6 73.4$\pm$0.4 75.2$\pm$0.6 74.8$\pm$0.6 75.8$\pm$2.6 68.6$\pm$1.4 75.7$\pm$0.8 75.2$\pm$3.0 68.4$\pm$1.1 77.2$\pm$0.7 77.3$\pm$7.3 63.5$\pm$1.1 70.5$\pm$2.4
Accuracy $\uparrow$ 79.1$\pm$0.7 72.0$\pm$0.7 72.4$\pm$0.5 73.6$\pm$0.6 73.2$\pm$0.4 68.3$\pm$2.6 77.6$\pm$0.9 72.7$\pm$1.2 66.8$\pm$3.0 77.6$\pm$1.0 71.8$\pm$1.2 68.4$\pm$6.8 78.0$\pm$0.6 73.5$\pm$1.4
DTO $\downarrow$ 43.6$\pm$0.6 38.4$\pm$0.5 38.3$\pm$0.4 36.2$\pm$0.1 36.7$\pm$0.4 39.9$\pm$3.6 38.6$\pm$0.7 36.6$\pm$0.5 41.4$\pm$4.1 38.8$\pm$0.6 36.2$\pm$1.2 40.0$\pm$3.5 42.6$\pm$0.8 39.7$\pm$1.8
FF-score $\uparrow$ 69.4$\pm$0.4 72.8$\pm$0.4 72.9$\pm$0.3 74.4$\pm$0.1 74.0$\pm$0.3 71.8$\pm$2.6 72.8$\pm$0.5 74.1$\pm$0.3 70.8$\pm$3.0 72.7$\pm$0.4 74.4$\pm$0.8 71.9$\pm$2.2 70.0$\pm$0.6 71.9$\pm$1.3
MOJI balanced Fairness $\uparrow$ 69.5$\pm$0.2 83.8$\pm$0.8 84.7$\pm$1.5 85.5$\pm$0.5 85.6$\pm$0.6 79.7$\pm$3.9 77.1$\pm$0.9 86.6$\pm$0.5 77.6$\pm$4.2 77.0$\pm$0.8 87.5$\pm$0.5 85.8$\pm$8.3 71.7$\pm$0.6 77.9$\pm$4.4
Accuracy $\uparrow$ 71.9$\pm$0.4 74.0$\pm$0.4 74.1$\pm$0.6 74.8$\pm$0.3 74.5$\pm$0.4 73.6$\pm$0.8 74.0$\pm$0.3 74.0$\pm$0.2 73.0$\pm$1.2 74.0$\pm$0.4 73.7$\pm$0.5 63.0$\pm$6.9 71.9$\pm$0.4 68.5$\pm$2.0
DTO $\downarrow$ 41.5$\pm$0.4 30.7$\pm$0.7 30.1$\pm$0.7 29.0$\pm$0.1 29.3$\pm$0.4 33.4$\pm$3.0 34.7$\pm$0.7 29.3$\pm$0.3 35.2$\pm$3.7 34.7$\pm$0.6 29.1$\pm$0.6 40.9$\pm$4.4 39.9$\pm$0.6 38.8$\pm$1.2
FF-score $\uparrow$ 70.7$\pm$0.3 78.6$\pm$0.5 79.1$\pm$0.6 79.8$\pm$0.1 79.6$\pm$0.3 76.5$\pm$2.2 75.5$\pm$0.5 79.8$\pm$0.2 75.2$\pm$2.6 75.5$\pm$0.4 80.0$\pm$0.4 71.9$\pm$3.2 71.8$\pm$0.4 72.8$\pm$1.0
BIOS-2 imbalanced Fairness $\uparrow$ 90.4$\pm$0.8 97.2$\pm$0.8 96.4$\pm$0.4 95.8$\pm$1.0 96.6$\pm$0.8 92.8$\pm$9.3 93.0$\pm$2.3 94.5$\pm$4.4 77.3$\pm$6.5 94.8$\pm$2.3 96.7$\pm$0.9 92.0$\pm$1.6 91.7$\pm$1.4 91.9$\pm$1.7
Accuracy $\uparrow$ 96.7$\pm$0.1 94.8$\pm$0.4 95.0$\pm$0.3 95.2$\pm$0.3 95.0$\pm$0.5 60.5$\pm$3.6 94.6$\pm$0.2 92.0$\pm$0.4 64.0$\pm$5.5 94.6$\pm$0.1 93.2$\pm$0.3 95.8$\pm$0.6 96.2$\pm$0.3 95.8$\pm$0.6
DTO $\downarrow$ 10.1$\pm$0.7 5.9$\pm$0.2 6.2$\pm$0.2 6.5$\pm$0.6 6.1$\pm$0.3 41.3$\pm$2.1 9.0$\pm$1.7 10.3$\pm$2.8 43.4$\pm$2.4 7.7$\pm$1.7 7.6$\pm$0.5 9.1$\pm$1.3 9.1$\pm$1.3 9.2$\pm$1.4
FF-score $\uparrow$ 93.5$\pm$0.4 96.0$\pm$0.2 95.7$\pm$0.1 95.5$\pm$0.4 95.8$\pm$0.2 72.8$\pm$2.3 93.8$\pm$1.2 93.2$\pm$2.3 69.6$\pm$1.7 94.7$\pm$1.2 94.9$\pm$0.4 93.8$\pm$0.7 93.9$\pm$0.7 93.8$\pm$0.8
BIOS-2 balanced Fairness $\uparrow$ 89.7$\pm$0.6 97.8$\pm$0.8 98.0$\pm$0.8 95.9$\pm$0.8 96.4$\pm$0.3 90.6$\pm$9.8 93.7$\pm$2.6 94.6$\pm$4.2 74.8$\pm$9.2 96.6$\pm$1.8 97.5$\pm$0.9 91.8$\pm$2.0 91.6$\pm$1.9 91.3$\pm$1.9
Accuracy $\uparrow$ 92.4$\pm$0.3 91.9$\pm$0.6 91.9$\pm$1.5 92.6$\pm$0.5 92.9$\pm$0.6 49.9$\pm$9.4 90.9$\pm$1.3 89.3$\pm$1.8 63.8$\pm$10.1 91.9$\pm$0.7 90.6$\pm$1.3 91.9$\pm$1.2 92.2$\pm$1.0 91.9$\pm$1.2
DTO $\downarrow$ 12.8$\pm$0.6 8.5$\pm$0.4 8.4$\pm$1.4 8.5$\pm$0.2 8.0$\pm$0.6 52.4$\pm$6.0 11.1$\pm$2.4 12.4$\pm$3.4 46.0$\pm$4.3 8.9$\pm$1.4 9.7$\pm$1.5 11.7$\pm$1.7 11.5$\pm$1.7 12.0$\pm$1.6
FF-score $\uparrow$ 91.1$\pm$0.4 94.7$\pm$0.1 94.9$\pm$0.7 94.2$\pm$0.2 94.6$\pm$0.3 63.0$\pm$4.6 92.3$\pm$1.9 91.9$\pm$2.9 67.5$\pm$3.0 94.2$\pm$1.2 93.9$\pm$1.1 91.8$\pm$1.1 91.9$\pm$1.2 91.6$\pm$1.1
@@
Table 1: Comparison of debiasing methods and selective debiasing. The best results in the group are in bold, and the best results overall are underlined. The results averaged over 5 random seeds.
@@
For the BIOS-2 dataset, we use BERT (“bert-base-cased”) Devlin et al. (2019). For the MOJI dataset, we use the domain-specific BERTweet model Nguyen et al. (2020) given that the data is from social media sources. For both models, we added a three-layer MLP as a classification head, following Han et al. (2022b). Model hyperparameters are described in Appendix A.
@@
We compare the proposed selective debiasing approach to inference-time debiasing of all predictions using LEACE and INLP, as well as to training-time debiasing techniques. We also compare the proposed KL-based bias quantification score with a UQ baseline: Softmax Response (SR: Geifman and El-Yaniv (2017)), calculated as $U_{SR}=1-\max_{c\in C}p_{c}$.

### v3

@@ Abstract
We propose selective debiasing – an inference-time safety mechanism designed to enhance the overall model quality in terms of prediction performance and fairness, especially in scenarios where retraining the model is impractical. The method draws inspiration from selective classification, where at inference time, predictions with low quality, as indicated by their uncertainty scores, are discarded. In our approach, we identify the potentially biased model predictions and, instead of discarding them, we remove bias from these predictions using LEACE – a post-processing debiasing method. To select problematic predictions, we propose a bias quantification approach based on KL divergence, which achieves better results than standard uncertainty quantification methods. Experiments on text classification datasets with encoder-based classification models demonstrate that selective debiasing helps to reduce the performance gap between post-processing methods and debiasing techniques from the at-training and pre-processing categories. [^1] The code is available online at https://github.com/glkuzi/selective-debiasing
Inference-Time Selective Debiasing to Enhance Fairness in
Text Classification Models
Gleb Kuzmin2,4   Neemesh Yadav5   Ivan Smirnov3,4 Timothy Baldwin1,6   Artem Shelmanov1 1MBZUAI    2Weakly-Supervised NLP Group    3RUDN University 4Laboratory for Analysis and Controllable Text Generation Technologies RAS 5IIIT Delhi    6The University of Melbourne kuzmin@airi.net    neemesh20529@iiitd.ac.in    ivs@isa.ru {timothy.baldwin, artem.shelmanov}@mbzuai.ac.ae
@@ 1 Introduction
Fairness is an important safety characteristic of a machine learning (ML) model, representing the model’s ability to classify instances without discrimination based on various sensitive attributes, such as race, gender, and age Blodgett et al. (2020). For the past few years, numerous works have investigated and promoted fairness, and a variety of fairness definitions have been proposed Blodgett et al. (2020); Han et al. (2022b). One prominent type of fairness is group fairness, also known as the equal opportunity criterion, which reflects the inequality of opportunities across different groups Han et al. (2022a). The inequality in the model predictions usually comes from inadequate or biased training data, and to address this problem and achieve better fairness, researchers have proposed various debiasing techniques Li et al. (2018); Han et al. (2021, 2022a); Belrose et al. (2023); Kuzmin et al. (2023). The majority of these techniques assume that one has access to the complete training data and the ability to retrain the model from scratch using some special loss function or reweighting the training instances. However, there are many situations when this assumption does not hold. There is a need for inference-time safety mechanisms that protect users from inadequate model behavior.
Inference-time safety mechanisms are primarily associated with uncertainty quantification (UQ) techniques Gal and Ghahramani (2016) and selective classification Geifman and El-Yaniv (2017); Xin et al. (2021); Vazhentsev et al. (2022, 2023). Selective classification aims to enhance the reliability of ML-based applications by abstaining from unreliable predictions with high uncertainty. We suggest that the same approach could be applied to increase fairness.
In this work, we propose an inference-time safety mechanism that aims to increase the overall quality of models in terms of prediction performance and fairness in situations when model retraining is prohibitive. We call this approach selective debiasing. Instead of rejecting predictions of selected instances as in selective classification, we apply to them inference-time debiasing using post-processing debiasing techniques. To the best of our knowledge, this style of approach is novel to the NLP community.
@@ 2 Background
Debiasing techniques can be categorized into three groups: at-training, pre-processing, and post-processing Han et al. (2022b).
@@
One of the most popular at-training methods is adversarial training (Adv) Li et al. (2018). It aims to solve a minimax game between minimizing the loss for the primary task and maximizing the loss for predicting the protected attribute. The diverse adversaries method (DAdv) Han et al. (2021) extends Adv by using an ensemble of multiple diverse discriminators instead of just one. In the pre-processing category, one of the most remarkable methods is Balanced Training with Equal Opportunity (BTEO) Han et al. (2022a). It rebalances the dataset to minimize the True Positive Rate (TPR) gap between two protected groups. In the same category, Balanced Training with Joint balance (BTJ) Lahoti et al. (2020) aims to improve the worst-case performance over all unobserved protected groups by focusing on the computationally identifiable regions of error.
@@
INLP is an iterative method that involves finding an orthogonal projection of a linear classifier matrix, which is initially learned to predict protected attributes from representations (e.g. hidden states of the standard model). This orthogonal projection is then iteratively used to remove all relevant information from these representations, which was used by the classifier to predict protected attributes.
LEACE is a concept erasure technique that renders representations impervious to the prediction of a specific concept while minimizing changes to the original representations. To construct a transformation matrix, it first whitens the data by equalizing the variance across all directions in the representation space. Next, the data is orthogonally projected onto the subspace that captures correlations between representations and protected attributes. Finally, the data is unwhitened using the same covariance matrix. This resulting transformation matrix is subtracted from the original representations (see the formal definition for LEACE in Appendix A).
At-training and pre-processing methods require retraining the model from scratch and access to the whole training set. They also cannot be selectively applied to a subset of predictions. Post-processing techniques do not involve changes to the model itself, can be trained on a subset of data, and can be applied to predictions selectively. However, their performance is usually worse.
In our work, we propose a method that combines the advantages of both post-processing and at-training / pre-processing methods. While it does not need access to the whole training dataset or retraining the model from scratch, it also has better performance than the standard post-processing techniques.
3 Proposed Method
We propose a selective approach, based on applying debiasing only to predictions with the highest bias score. This section introduces the general concept of selective debiasing and presents the bias quantification method underlying this approach.
@@ 3 Proposed Method
Selective classification is a widely recognized safety mechanism that safeguards against using unreliable model predictions. In this approach, predictions flagged as unreliable due to high uncertainty scores are handled differently, e.g. they are rejected or are escalated to human operators for further review.
Instead of rejecting instances completely as in selective classification, we apply debiasing to selected predictions. In particular, we identify the potentially most biased instances using a bias quantification method $\mathcal{B}(x_{i},p_{i})$ and replace the original prediction $p_{i}=f(x_{i})$ with a prediction debiased using a post-processing method $d$: $\hat{p_{i}}=d(f(x_{i}))$:
$\bar{p_{i}}=\begin{cases}p_{i}=f(x_{i}),\text{ if }\mathcal{B}(x_{i},p_{i})<h\\ \hat{p_{i}}=d(f(x_{i})),\text{ if }\mathcal{B}(x_{i},p_{i})\geq h,\end{cases}$ (1)
where $h$ is a predefined threshold selected on a validation set.
We note that the proposed approach is different from the standard post-processing debiasing methods since we change predictions for only some instances. While debiasing all predictions might significantly reduce model performance, modifying only predictions likely to be of low quality or biased is less risky in terms of worsening outcomes and has the potential to correct errors. Such an approach also allows tuning the accuracy–fairness trade-off for debiasing methods Han et al. (2022b); Kuzmin et al. (2023).
@@
Selective classification is usually based on UQ methods. However, uncertainty on its own does not reflect the presence of bias; it simply highlights potentially erroneous predictions. Figure 1 presents a motivational example. It shows the rejection plots for oracle rejection strategies in selective classification for both accuracy and fairness (see the exact definition of fairness in Appendix E). We can see that the fairness oracle outperforms the UQ oracle in terms of fairness while keeping the same performance in terms of accuracy. These results illustrate that it is possible to improve fairness without penalty to accuracy by changing the order of instances being eliminated, i.e. using a different selection criterion.
Consider a multi-label classification model with classes $c\in C$. To quantify how biased a model prediction is for a given instance, we suggest using the Kullback-Leibler (KL) divergence Kullback and Leibler (1951) between the originally predicted probability distribution $p_{i}^{c}$ and distribution $\hat{p_{i}}^{c}$ after debiasing:
$\mathcal{B}_{KL}^{i}=\sum_{c\in C}p_{i}^{c}\log\left(\frac{p_{i}^{c}}{\hat{p_{i}}^{c}}\right).$ (2)
KL divergence measures the difference in predictions between the standard and the debiased model. The greater the difference, the more information about the protected attribute is removed from the original representation of the instance. This approach could be used with various post-processing methods. In particular, we suggest using LEACE, but also present results with INLP.
Note that applying a post-processing method to a model is a matter of one or two matrix multiplications. An additional prediction step requires inferring only the last layer of a model, which is very fast. Therefore, the runtime overhead introduced by bias quantification is very small (see Appendix H).
@@ 4 Experiments
Figure 1: Rejection results for fairness and accuracy with oracle scores on a synthetic dataset with a LogReg model; the FR-AUC and Acc-AUC are the areas under fairness– and accuracy–rejection curves correspondingly. The details are presented in Appendix B.
@@ 4.1 Experimental Setup
For our experiments, we use two English text classification datasets that, in addition to target variables, provide explicit protected attributes. The first is MOJI Blodgett et al. (2016), a dataset for sentiment analysis with a binary class (“happy” and “sad”) and a binary protected attribute, which corresponds to the author’s ethnicity (African American English (AAE) vs. Standard American English (SAE)). The second is a version of the widely used BIOS dataset De-Arteaga et al. (2019) for occupation classification with a binary gender as the protected attribute. BIOS-2 Subramanian et al. (2021) is a two-class subsample of the original BIOS dataset with a highly-skewed joint distribution of classes and protected attribute values. As it has been shown to be beneficial to report results for both “balanced” and “imbalanced” versions of datasets Kuzmin et al. (2023), we conduct experiments on both versions. Detailed information and statistics of the datasets are presented in Appendix C. Due to the limited availability of datasets with annotated protected attributes, most research on debiasing and fairness has been conducted on these few datasets Han et al. (2022b).
Debiasing method type No debiasing At-training Pre-processing Post-processing & Selective
@@
MOJI imbalanced Fairness $\uparrow$ 61.8$\pm$0.7 73.7$\pm$0.6 73.4$\pm$0.4 75.2$\pm$0.6 74.8$\pm$0.6 75.8$\pm$2.6 68.6$\pm$1.4 75.7$\pm$0.8 75.2$\pm$3.0 68.4$\pm$1.1 77.2$\pm$0.7
Accuracy $\uparrow$ 79.1$\pm$0.7 72.0$\pm$0.7 72.4$\pm$0.5 73.6$\pm$0.6 73.2$\pm$0.4 68.3$\pm$2.6 77.6$\pm$0.9 72.7$\pm$1.2 66.8$\pm$3.0 77.6$\pm$1.0 71.8$\pm$1.2
DTO $\downarrow$ 43.6$\pm$0.6 38.4$\pm$0.5 38.3$\pm$0.4 36.2$\pm$0.1 36.7$\pm$0.4 39.9$\pm$3.6 38.6$\pm$0.7 36.6$\pm$0.5 41.4$\pm$4.1 38.8$\pm$0.6 36.2$\pm$1.2
FF-score $\uparrow$ 69.4$\pm$0.4 72.8$\pm$0.4 72.9$\pm$0.3 74.4$\pm$0.1 74.0$\pm$0.3 71.8$\pm$2.6 72.8$\pm$0.5 74.1$\pm$0.3 70.8$\pm$3.0 72.7$\pm$0.4 74.4$\pm$0.8
MOJI balanced Fairness $\uparrow$ 69.5$\pm$0.2 83.8$\pm$0.8 84.7$\pm$1.5 85.5$\pm$0.5 85.6$\pm$0.6 79.7$\pm$3.9 77.1$\pm$0.9 86.6$\pm$0.5 77.6$\pm$4.2 77.0$\pm$0.8 87.5$\pm$0.5
Accuracy $\uparrow$ 71.9$\pm$0.4 74.0$\pm$0.4 74.1$\pm$0.6 74.8$\pm$0.3 74.5$\pm$0.4 73.6$\pm$0.8 74.0$\pm$0.3 74.0$\pm$0.2 73.0$\pm$1.2 74.0$\pm$0.4 73.7$\pm$0.5
DTO $\downarrow$ 41.5$\pm$0.4 30.7$\pm$0.7 30.1$\pm$0.7 29.0$\pm$0.1 29.3$\pm$0.4 33.4$\pm$3.0 34.7$\pm$0.7 29.3$\pm$0.3 35.2$\pm$3.7 34.7$\pm$0.6 29.1$\pm$0.6
FF-score $\uparrow$ 70.7$\pm$0.3 78.6$\pm$0.5 79.1$\pm$0.6 79.8$\pm$0.1 79.6$\pm$0.3 76.5$\pm$2.2 75.5$\pm$0.5 79.8$\pm$0.2 75.2$\pm$2.6 75.5$\pm$0.4 80.0$\pm$0.4
BIOS-2 imbalanced Fairness $\uparrow$ 90.4$\pm$0.8 97.2$\pm$0.8 96.4$\pm$0.4 95.8$\pm$1.0 96.6$\pm$0.8 92.8$\pm$9.3 93.0$\pm$2.3 94.5$\pm$4.4 77.3$\pm$6.5 94.8$\pm$2.3 96.7$\pm$0.9
Accuracy $\uparrow$ 96.7$\pm$0.1 94.8$\pm$0.4 95.0$\pm$0.3 95.2$\pm$0.3 95.0$\pm$0.5 60.5$\pm$3.6 94.6$\pm$0.2 92.0$\pm$0.4 64.0$\pm$5.5 94.6$\pm$0.1 93.2$\pm$0.3
DTO $\downarrow$ 10.1$\pm$0.7 5.9$\pm$0.2 6.2$\pm$0.2 6.5$\pm$0.6 6.1$\pm$0.3 41.3$\pm$2.1 9.0$\pm$1.7 10.3$\pm$2.8 43.4$\pm$2.4 7.7$\pm$1.7 7.6$\pm$0.5
FF-score $\uparrow$ 93.5$\pm$0.4 96.0$\pm$0.2 95.7$\pm$0.1 95.5$\pm$0.4 95.8$\pm$0.2 72.8$\pm$2.3 93.8$\pm$1.2 93.2$\pm$2.3 69.6$\pm$1.7 94.7$\pm$1.2 94.9$\pm$0.4
BIOS-2 balanced Fairness $\uparrow$ 89.7$\pm$0.6 97.8$\pm$0.8 98.0$\pm$0.8 95.9$\pm$0.8 96.4$\pm$0.3 90.6$\pm$9.8 93.7$\pm$2.6 94.6$\pm$4.2 74.8$\pm$9.2 96.6$\pm$1.8 97.5$\pm$0.9
Accuracy $\uparrow$ 92.4$\pm$0.3 91.9$\pm$0.6 91.9$\pm$1.5 92.6$\pm$0.5 92.9$\pm$0.6 49.9$\pm$9.4 90.9$\pm$1.3 89.3$\pm$1.8 63.8$\pm$10.1 91.9$\pm$0.7 90.6$\pm$1.3
DTO $\downarrow$ 12.8$\pm$0.6 8.5$\pm$0.4 8.4$\pm$1.4 8.5$\pm$0.2 8.0$\pm$0.6 52.4$\pm$6.0 11.1$\pm$2.4 12.4$\pm$3.4 46.0$\pm$4.3 8.9$\pm$1.4 9.7$\pm$1.5
FF-score $\uparrow$ 91.1$\pm$0.4 94.7$\pm$0.1 94.9$\pm$0.7 94.2$\pm$0.2 94.6$\pm$0.3 63.0$\pm$4.6 92.3$\pm$1.9 91.9$\pm$2.9 67.5$\pm$3.0 94.2$\pm$1.2 93.9$\pm$1.1
@@
Table 1: Comparison of debiasing methods and selective debiasing. The best results in the group are in bold, and the best results overall are underlined. The results are averaged over 5 random seeds. The gray color corresponds to the results with p-value > 0.05 with respect to standard model.
Metrics.
We employ several metrics to evaluate the predictive performance and fairness of the model. To evaluate the performance, we use accuracy. For fairness, we consider the widely used equal opportunity criterion Hardt et al. (2016); Han et al. (2022a, b). We also use two aggregated metrics to evaluate the performance in terms of both accuracy and fairness. The first one is the distance to the optimal point (DTO) Han et al. (2021):
$\scriptstyle\text{DTO}=\sqrt{(1-\text{Accuracy})^{2}+(1-\text{Fairness})^{2}}.$ (3)
[…]
@@
For the BIOS-2 dataset, we use BERT (“bert-base-cased”) Devlin et al. (2019). For the MOJI dataset, we use the domain-specific BERTweet model Nguyen et al. (2020) which is good for processing data from social media sources. For both models, we add a three-layer MLP as a classification head, following Han et al. (2022b). Model hyperparameters are described in Appendix D.
@@
We compare the proposed selective debiasing approach to inference-time debiasing of all predictions using LEACE and INLP, as well as to at-training and pre-processing debiasing techniques: Adv, DAdv, BTEO, BTJ. We also compare the proposed KL-based bias quantification score with a UQ baseline: Softmax Response (SR: Geifman and El-Yaniv (2017)), calculated as $\mathcal{B}_{SR}(x_{i})=1-\max_{c\in C}p^{c}_{i}$.
Details of debiasing methods.
Pre-processing and at-training debiasing methods were applied while training the model from scratch on the full dataset, whereas post-processing methods were trained using only 20% of the data. The optimal threshold for selective debiasing was chosen based on the first 15% of the validation set. “LEACE-last” in our experiments represents LEACE applied to the outputs of the last hidden layer of the classifier, while “LEACE-cls” is LEACE applied to each linear layer of the classification head of the model. The hyperparameters of debiasing methods are provided in Appendix D.